X+2y=11 over 3x + 4y =23 solve for x and y
Answer by ewatrrr(24785)

 
Hi,
x + 2y = 11
3x + 4y = 23
-3x -6y = 11*-3 |multiplying 1st EQ thru by -3 to eliminate x
3x +4y = 23
-2y = -10
y = 5 and x = 1 (x = -2*5+11)
Ordered pair (1,5) the solution for this system of equations
CHECKING our Answer***
3*1 + 4*5 = 23